1. The following transactions are July activities of Josh’s video service:

-Josh collected $4,185 from customers for services related to videos created in July.

-Josh billed a customer for $811 for a video made on the last day of July. The bill is to be paid in August.

-On July 1, a customer gave Josh advance payments totaling $3,382 for the four months of work to be done equally staring in July

-Josh received $894 from credit sales made to customers last month (in June).

By what amount did total assets increase during July?

2.

ABC Company had the following account balances.

Accounts Payable $ 948
Accounts Receivable $ 1,694
Cash $ 4,736
Common Stock $ 1,374
Deferred Revenue $ 1,007

ABC Company was missing the account balance for retained earnings. What is ABC Company's balance in retained earnings?

3.

During its first month of operations, ABC Company provided $96,229 of services to their customers. ABC Company’s costs included rent of $18,632 and salaries and wages totaling $24,737. At the end of the month, the board of directors approved a dividend of $2,637. What is the Retained Earnings for ABC Company at the end of the month?

4.

The following transactions are July activities of Josh's video service:

  1. Josh paid $575 to plumbers for repairing a broken pipe in the restrooms.
  2. Josh paid $461 for the June electricity bill and received the July bill for $716, which will be paid in August.
  3. Josh paid 231 to employees for work in June.

What are the total change to liabilities that Noah will recognize in July?

5. The following transactions are July activities of Josh's video service:

  • Josh collected $1,321 from credit customers for services related to videos created in June.        
  • Josh billed a customer for $849 for a video made on the last day of July. The bill is to be paid in August.                               
  • On July 1, a customer gave Josh advance payments totaling $1,600 for the five months of work to be done equally staring in July
  • Josh received $788 from video created for customers in July.

What is the revenue Josh will recognize in July?
Answer #1

Answer-1:

Cash = $4,185 + 3,382 + 894 =8,461

Account receivable = $811 - 894 = -83

Total assets increase during July = $8,461 - 83 = $8,378

Answer-2:

ABC Company's balance in retained earnings = $3,101

Explanation:

Assets Account receivable Cash $ Liabilities & Equity 1,694 Account payable 4,736 Common stock Deferred revenue Retained earn

Answer-3:

The Retained Earnings for ABC Company at the end of the month = $50,223

Explanation:

$ Service Revenue Less: Rent expense Less: Salaries and wages Net Income 96,229 18,632 24,737 52,860 Beginning Retained earni

Answer-4:

The total change to liabilities that Noah will recognize in July = -$461 + 716 - 231 = $24 increase

Answer-5:

The revenue Josh will recognize in July = $849 + (1,600 / 5) + 788 = $1,957